# Holographic duality from Howe duality: Chern–Simons gravity as an ensemble of code CFTs

### v1

@@ 4 Discussion
This paper has two main goals. The first is to unify recent literature about ensembles of CFTs based on codes over ${\mathbb{F}}_{p}\times{\mathbb{F}}_{p}$ of length $n$. We reviewed the construction of Chern–Simons theory behind these ensembles and explained that each individual code CFT is holographically dual to a trivial TQFT in the bulk emerging after gauging of a particular maximal non-anomalous subgroup of the one-form symmetry, as outlined in Barbar:2023ncl ; Benini:2022hzx ; Aharony:2023zit . We have also shown that the Hilbert space ${\mathcal{H}}^{g}$ of the Chern–Simons theory on a genus $g$ Riemann surface is equal to the tensor product of $g$ Hilbert spaces of $2n$ level-$p$ qudits. Gauging a maximal subgroup of one-form symmetry projects the Hilbert space to a one-dimensional “code subspace” – which, in terms of the qubits, is defined by a CSS self-dual quantum stabilizer code. In this language, a state of the Chern–Simons theory on any particular handlebody is also a quantum stabilizer state, providing a quantum information theory interpretation of the holographic duality (1).
@@ 4.1 Genus reduction and the sum over all topologies
First we would like to place our results in the context of Dymarsky:2024frx , which has shown that any 3d TQFT summed over all possible topologies is holographically dual to a probabilistic ensemble of boundary CFTs, defined by maximal gaugings of the bulk theory. The proof there is rigorous for any abelian Chern–Simons theory, although there are many subtleties for more complicated theories e.g. those with a continuous spectrum, such as Virasoro TQFT Collier:2023fwi ; Collier:2024mgv . The general picture is that when the boundary Riemann surface has a very large genus $g$, the holographic duality trivializes and only handlebodies contribute to the bulk sum. All other non-trivial topologies then emerge via genus reduction. An important point is that genus reduction unavoidably yields singular topologies, which can be understood as handlebodies with Wilson line insertions.
What our paper shows is that in certain simpler cases, such as the case of prime $p$ of this paper, the contribution of all possible topologies, including the singular ones, is proportional to a sum that only includes handlebodies, and hence merely renormalizes the overall coefficient. This statement applies to the sum, not to individual topologies. The genus reduction can be formulated in quantum mechanical terms, as a particular linear map of a state from ${\mathcal{H}}^{g}$ to ${\mathcal{H}}^{g-1}$, see Dymarsky:2024frx . What our results imply, and which would be difficult to see directly, is that when $p$ is a prime the genus reduction maps the sum over the handlebodies of genus $g$ to the sum over the handlebodies of genus $g-1$.
@@ 4.2 $U(1)$ gravity
An interesting question is whether our ensemble over a discrete set of lattice theories gives rise to the average over the whole Narain moduli space dual to “$U(1)$-gravity” of Afkhami-Jeddi:2020ezh ; Maloney:2020nni in the limit $p\rightarrow\infty$. [^27] With the exception of the case of $n=1$. Strictly speaking at this point the relation to the average over the whole moduli space is conjectural; although in this limit the number of code CFTs becomes infinite, it is an open question to show that they will densely and homogeneously – with respect to Zamolodchikov metric – populate the Narain moduli space in the large-$p$ limit. Evidence in support of the limit was given in Aharony:2023zit , which evaluated the code CFT averaged torus partition function to become the Eisenstein series (3). A related check can be found in Appendix G, where we show that the gap of $U(1)$ primaries in the large-$p$ limit approaches $\Delta^{*}=1/(2\pi e)$ expected for $U(1)$-gravity Afkhami-Jeddi:2020ezh .
@@ 4.3 Howe duality for other fields
The theta correspondence can be thought of as a specific example of a more general principle – indeed, there are examples in physics with two groups acting on a physical system and exhibiting a special relationship among their representations Howe:1987tv ; Rowe:2012ym ; Basile:2020gqi . Here typically one of the groups is a symmetry group, and another is a “dynamical group.” Representations of the two groups are in one-to-one correspondence and can be indexed by a joint label. A concrete example is the central potential problem, e.g. the hydrogen atom. In the latter case a basis of eigenfunctions can be taken to be $\psi_{nlm}$, where $(n,l)$ are labels for the radial problem furnishing irreps of the dynamical group $SU(1,1)$, and $(l,m)$ are labels for the spherical problem, furnishing irreps of the symmetry group $SO(3)$. The common label (corresponding to $i$ in (104)) is $l$, which determines the rotational irrep with $12l+$ 1 states labeled by $m$, but also a highest-weight irrep of $SU(1,1)$, with infinitely many states labeled by $n\geqslant l+1$ Rowe:2012ym . More examples of this kind can be found in Rowe:2012ym ; Basile:2020gqi , demonstrating that the dual pair correspondence is common in physics.
@@
For global fields, including ${\mathbb{Q}}$ and its finite extensions, Howe duality manifests in the theory of automorphic forms – functions on spaces defined by a group $G$ with certain transformation properties under a subgroup $\Gamma\subset G$, plus other conditions on smoothness/fall-off. conditions. Howe duality becomes a correspondence between the automorphic forms of the two groups in the dual pair, which takes the form of a map between automorphic representations of the two groups. A general treatment requires introducing the so-called adeles ${\mathbb{A}}$ of a local field $F$ [^28] This is the (restricted) direct product of all completions of $F$ with respect to each of its valuations. If $F={\mathbb{Q}}$, then an element of ${\mathbb{A}}={\mathbb{A}}_{\mathbb{Q}}$ is an infinite tuple consisting of elements in ${\mathbb{R}}$, ${\mathbb{Q}}_{2}$, ${\mathbb{Q}}_{3}$, ${\mathbb{Q}}_{5}$, …where ${\mathbb{Q}}_{p}$ are the $p$-adics. There is an additional restriction that all but a finite number of the $p$-adic elements are integers. The motivation of introducing the adeles is that the rationals inside the adeles form a discrete subset, just like the integers inside ${\mathbb{R}}$. Statements including the familiar cosets $G({\mathbb{Z}})\backslash G({\mathbb{R}})$ can then be rephrased as statements for $G({\mathbb{Q}})\backslash G({\mathbb{A}})$, where the advantage is now that ${\mathbb{Q}}$ is a field, with nicer algebraic properties than the ring ${\mathbb{Z}}$. and the structure of a pairing between representations is formulated as a “theta lift,” which relates automorphic forms of the two commuting subgroups.
@@
[…]
@@
The relation (139) was first derived by Siegel Siegel1935 , generalizing the earlier Smith–Minkowski–Siegel mass formula Smith1868 ; Minkowski1885 . It was conjectured in Dymarsky:2020qom , also see Henriksson:2022dml , relation (139) can be interpreted in CFT terms as an average over the ensemble of chiral code-based lattice CFTs dual to chiral Chern-Simons theory. Note however that full set of chiral (meromorphic) CFT of central charge $n=0$ (mod 8) includes other theories (see Schellekens:1992db ; Lam2019 for $n=24$), and it is not obvious how to perform the average over this whole set.
@@ 4.4 Other code constructions
In fact, a preliminary study shows that whenever $k$ is square-root free, the orthogonal group acts transitively on codes over ${\mathbb{Z}}_{k}\times{\mathbb{Z}}_{k}$ and there is still a unique vector in ${\mathcal{H}}^{g}$ invariant under both groups. At the same time, providing a rigorous proof in that case is a non-trivial task, in part because ${\mathbb{Z}}_{k}$ is no longer a field.
@@
The mathematical statements of finite-field Howe duality apply for general fields ${\mathbb{F}}_{q}$, where $q=p^{n}$, and it is possible to consider averaging over ensembles of such codes. However, to provide an interpretation of this result in terms of CFTs or Chern–Simons theory would require new code-based CFT constructions. Note that ${\mathbb{F}}_{p^{n}}$ is not the same as the ring ${\mathbb{Z}}_{p^{n}}$ – the latter emerges from the construction of this paper with $k=p^{n}$, but there is no known mathematical statement similar to finite-field Howe duality in this case.
@@ 4.7 Holography and quantum information
As a byproduct of formulating the statement of holographic duality (1) in terms of group representation theory, we have arrived at the formulation in terms of families of quantum stablizer states, given by (103). This formulation provides an intriguing interpretation to (1) in terms of quantum information theory, and resonates with the lore that quantum codes of various kind must play a crucial role in holography Almheiri:2014lwa ; Pastawski:2015qua . It is given that quantum stabilizer codes, that are based on the commutative (abelian) groups $S$, is the artifact of the abelian nature of fusion rules of the bulk CS theory. More complicated, and interesting models of TQFT gravity would call for a generalization of this simple construction. It nevertheless would be very interesting to generalize the approach of quantum codes to describe the holographic duality in more general setting.
@@ Appendix A The Heisenberg group and the oscillator representation
We shall start by considering the field ${\mathbb{R}}$, which is simpler and more commonly considered, and then we will explain the situation for finite fields. Consider a quantum particle, free or with some potential, in $N$ spatial dimensions. Then it is described by position and momentum operators which satisfy canonical commutation relations of the form
[…]
The elements $\left\{P_{1},\ldots,P_{N},Q_{1},\ldots Q_{N},C\right\}$, where the central element $C$ will be chosen to be $-i$, form the Heisenberg Lie algebra $\mathfrak{h}_{2N+1}$. The algebra elements can be identified with $(N+2)\times(N+2)$ nilpotent matrices. For example, for $N=1$, we have
@@
so a general element given by $h=a\,P+b\,Q+c\,C$ is identified with
@@
When exponentiated, $Q$ acts on the wavefunctions by multiplying by a coordinate-dependent phase and $P$ generates a translation. This gives rise to the Schrödinger representation of the Heisenberg group. Considering $N$ dimensions, a general element $\pi_{S}\in H_{N+2}$ acts on the wavefunction $\psi(\textbf{q})\in L^{2}(\bf q)$ by
@@
$\left[\sum_{i}\left(a_{i}P_{i}+b_{i}Q_{i}\right),\sum_{j}\left(a^{\prime}_{j}P_{j}+b^{\prime}_{j}Q_{j}\right)\right]=-iS((\mathbf{a},\mathbf{b}),(\mathbf{a}^{\prime},\mathbf{b}^{\prime})).$ (151)
@@
and satisfy $[a_{i},a_{j}^{\dagger}]\ =\ \delta_{ij}$. With these we can define the infinite dimensional oscillator representation $\omega$, which is the minimal faithful representation of the metaplectic group. The oscillator representation is
@@ A.1 Finite fields
The finite-field Heisenberg group also obeys a version of the Stone–von Neumann theorem. For any group $G$ and rep $\rho$, the restriction of the representation to the center $\rho(Z(G))$ must be one-dimensional by Schur’s lemma. Therefore in an irreducible representation, $z\in Z(G)$ acts by multiplication by a (complex) scalar $\chi(z)$. The function $\chi$ is called the central character of $\rho$. In the case of the Heisenberg group, the center is $(\mathbf{0},\mathbf{0},c)\simeq{\mathbb{F}}_{p}$. The statement of uniqueness is that the representations of the Heisenberg group are uniquely (up to unitary equivalences) determined by the value of the central elements in the representation, i.e. by the choice of central character. There are $p$ central characters $\chi(z)=\xi^{z}$, where $\xi$ is a $p^{\text{th}}$ root of unity.
@@
Now $U_{g}$ forms a representation of the symplectic group with dimension $p^{N}$– this is known as the oscillator representation. It is naturally a projective representation, but (unlike in the case for local fields) the phases can be chosen so that it is a regular rep. We make no distinction between the two (projective / regular) representations in our notation, though they differ by a phase in the definition of $U_{g}$.
@@
The space of dimension $p^{N}$ (here, the space of functions) which carries a representation of the Heisenberg and symplectic groups is what we call the oscillator representation, denoted $\mathcal{X}$ in our paper. For us, it is spanned by the set $\psi_{v_{1},...,v_{g}}$ for codewords $v_{i}$ of length $2n$,
@@
The oscillator representation $\mathcal{X}$ is an irrep of the Heisenberg group, but actually breaks into two irreps of the symplectic group. On the space of functions, these reps are the even and odd functions, which are preserved by the symplectic group. These subspaces have dimensions $(p^{N}+1)/2$ and $(p^{N}-1)/2$, respectively and correspond to charge-conjugation even (odd). We can see this at work in the case where $p=3$, $n=1$, $g=1$, presented in section 3.3. There we find that the oscillator representation is $93^{2}=$ 9 dimensional, and it breaks into symplectic irreps of dimension $5$ and $4$. More generally, the group $Sp(2N,{\mathbb{F}}_{q})$ has, apart from the trivial representation, four “minimal” representations with dimensions Gurevich2016
@@ Appendix B Classical groups over finite fields
where ϵ=±italic-$\epsilon=\pm$ is the Witt index of the form $\eta$. We are interested in, c.f. (20),
@@ Appendix C Phase ambiguity for $\boldsymbol{p=3}$ (mod $\boldsymbol{4}$)
where $f(x)\in L^{2}({\mathbb{F}}_{p}^{2ng})$. Now while the first of these expressions has a more natural action, it is the second that is compatible with the oscillator representation being a regular representation of $Sp(4ng,{\mathbb{F}}_{p})$, see Neuhauser2002 . In other words, if we have $G\times H=O^{+}(2n,{\mathbb{F}}_{p})\times Sp(2g,{\mathbb{F}}_{p})$ as a subgroup of $\mathcal{S}=Sp(4gn,{\mathbb{F}}_{p})$, and $\omega$ is the oscillator representation, we then have the restrictions
@@
This extra phase factor $(\det\mathsf{g})^{\frac{p-1}{2}g}$ (the tensor product with a power of the determinant rep) is sometimes not made explicit in the literature, see however equation (1.4) of Ma2022 . [^35] We thank Shu-Yen Pan for correspondence. It also introduces an ambiguity in the nomenclature of the singlet representation: the case we need for our proof in the main text is where $\pi$ acts trivially, $\pi(\mathsf{g})=\mathds{1}$; for odd $\frac{p-1}{2}g$ this would correspond to the case where $U_{\mathsf{g}}$ acts as the determinant representation $U_{\mathsf{g}}=\det(\mathsf{g})\mathds{1}$.

### v2

@@ 4 Discussion
This paper has two main goals. The first is to unify recent literature about ensembles of CFTs based on codes over ${\mathbb{F}}_{p}\times{\mathbb{F}}_{p}$ of length $n$. We reviewed the construction of Chern–Simons theory behind these ensembles and explained that each individual code CFT is holographically dual to a trivial TQFT in the bulk emerging after gauging of a particular maximal non-anomalous subgroup of the one-form symmetry, as outlined in Barbar:2023ncl ; Benini:2022hzx ; Aharony:2023zit . We have also shown that the Hilbert space ${\mathcal{H}}^{g}$ of Chern–Simons theory on a genus-$g$ Riemann surface is equal to the tensor product of $g$ Hilbert spaces of $2n$ level-$p$ qudits. Gauging a maximal subgroup of one-form symmetry projects the Hilbert space to a one-dimensional “code subspace” – which, in terms of the qubits, is defined by a CSS self-dual quantum stabilizer code. In this language, a state of the Chern–Simons theory on any particular handlebody is also a quantum stabilizer state, providing a quantum information theory interpretation of the holographic duality (1).
@@ 4.1 Genus reduction and the sum over all topologies
First we would like to place our results in the context of Dymarsky:2024frx , which has shown that any 3d TQFT summed over all possible topologies is holographically dual to a probabilistic ensemble of boundary CFTs, defined by maximal gaugings of the bulk theory. The proof there is rigorous for any abelian Chern–Simons theory, although there are many subtleties for more complicated theories e.g. those with a continuous spectrum, such as Virasoro TQFT Collier:2023fwi ; Collier:2024mgv ; Takahashi:2024ukk . The general picture is that when the boundary Riemann surface has a very large genus $g$, the holographic duality trivializes and only handlebodies contribute to the bulk sum. All other non-trivial topologies then emerge via genus reduction. An important point is that genus reduction unavoidably yields singular topologies, which can be understood as handlebodies with Wilson line insertions.
What our paper shows is that in certain simpler cases, such as the case of prime $p$ of this paper, the contribution of all possible topologies, including the singular ones, is proportional to a sum that only includes handlebodies, and hence merely renormalizes the overall coefficient. This statement applies to the sum, not to individual topologies. The genus reduction can be formulated in quantum mechanical terms, as a particular linear map of a state from ${\mathcal{H}}^{g}$ to ${\mathcal{H}}^{g-1}$, see Dymarsky:2024frx . What our results imply, and which would be difficult to see directly, is that when $p$ is a prime the genus reduction maps the sum over the handlebodies of genus $g$ to the sum over the handlebodies of genus g1g-1.
@@ 4.2 $U(1)$ gravity
An interesting question is whether our ensemble over a discrete set of lattice theories gives rise to the average over the whole Narain moduli space dual to “$U(1)$-gravity” of Afkhami-Jeddi:2020ezh ; Maloney:2020nni in the limit $p\rightarrow\infty$. [^29] With the exception of the case of $n=1$. Strictly speaking, at this point the relation to the average over the whole moduli space is conjectural. Although it is known that in the large-$p$ limit the number of code CFTs becomes infinite, it is an open question to show that they will densely and homogeneously – with respect to Zamolodchikov metric – populate the Narain moduli space. Evidence in support of the limit was given in Aharony:2023zit , which evaluated the code CFT averaged torus partition function to become the Eisenstein series (3). A related check can be found in appendix G, where we show that the gap of $U(1)$ primaries in the large-$p$ limit approaches $\Delta^{*}=1/(2\pi e)$ expected for $U(1)$-gravity Afkhami-Jeddi:2020ezh .
@@ 4.3 Howe duality for other fields
The theta correspondence can be thought of as a specific example of a more general principle – indeed, there are examples in physics with two groups acting on a physical system and exhibiting a special relationship among their representations Howe:1987tv ; Rowe:2012ym ; Basile:2020gqi . Here typically one of the groups is a symmetry group, and another is a “dynamical group.” Representations of the two groups are in one-to-one correspondence and can be indexed by a joint label. A concrete example is the central potential problem, e.g. the hydrogen atom. In this case a basis of eigenfunctions can be taken to be $\psi_{nlm}$, where $(n,l)$ are labels for the radial problem furnishing irreps of the dynamical group $SU(1,1)$, and $(l,m)$ are labels for the spherical problem, furnishing irreps of the symmetry group $SO(3)$. The common label (corresponding to $i$ in (104)) is $l$, which determines the rotational irrep with $2l+1$ states labeled by $m$, but also a highest-weight irrep of $SU(1,1)$, with infinitely many states labeled by $n\geqslant l+1$ Rowe:2012ym . More examples of this kind can be found in Rowe:2012ym ; Basile:2020gqi , demonstrating that the dual pair correspondence is common in physics.
@@
[…]
@@
$\int d\mu(m)\,\Theta_{\Lambda}(\Omega,m)=E_{\frac{n}{2}}(\Omega)/(\det\operatorname*{Im}\Omega)^{n},$ (138)
@@
The relation (139) was first derived by Siegel Siegel1935 , generalizing the earlier Smith–Minkowski–Siegel mass formula Smith1868 ; Minkowski1885 . It was conjectured in Dymarsky:2020qom , also see Henriksson:2022dml , that the relation (139) can be interpreted in CFT terms as an average over the ensemble of chiral code-based lattice CFTs dual to chiral Chern–Simons theory. Note however that full set of chiral (meromorphic) CFT of central charge $n=0$ (mod 8) includes other theories (see Schellekens:1992db ; Lam2019 for $n=24$), and it is not obvious how to perform the average over this whole set.
@@ 4.4 Other code constructions
In fact, a preliminary study shows that whenever $k$ is square-free, the orthogonal group acts transitively on codes over ${\mathbb{Z}}_{k}\times{\mathbb{Z}}_{k}$ and there is still a unique vector in ${\mathcal{H}}^{g}$ invariant under both groups. At the same time, providing a rigorous proof in that case is a non-trivial task, in part because ${\mathbb{Z}}_{k}$ is no longer a field.
@@
The mathematical statements of finite-field Howe duality apply for general fields ${\mathbb{F}}_{q}$, where $q=p^{m}$, and it is possible to consider averaging over ensembles of such codes with $m>1$. However, to provide an interpretation of this result in terms of CFTs or Chern–Simons theory would require new code-based CFT constructions. Note that ${\mathbb{F}}_{p^{m}}$ is not the same as the ring ${\mathbb{Z}}_{p^{m}}$ – the latter emerges from the construction of this paper with $k=p^{m}$, but there is no known mathematical statement similar to finite-field Howe duality in this case.
@@ 4.7 Holography and quantum information
As a byproduct of formulating the statement of holographic duality (1) in terms of group representation theory, we arrived at the formulation in terms of families of quantum stablizer states, given by (103). This formulation provides an intriguing interpretation to (1) in terms of quantum information theory, and resonates with the lore that quantum codes of various kind must play a crucial role in holography Almheiri:2014lwa ; Pastawski:2015qua . It is given that quantum stabilizer codes, that are based on the commutative (abelian) groups $S$, is the artifact of the abelian nature of fusion rules of the bulk CS theory. More complicated, and interesting models of TQFT gravity would call for a generalization of this simple construction. It nevertheless would be very interesting to generalize the approach of quantum codes to describe the holographic duality in more general setting.
@@ Appendix A The Heisenberg group and the oscillator representation
We shall start by considering the field $\mathbb{R}$}, which is simpler and more commonly considered, and then we will explain the situation for finite fields. Consider a quantum particle, free or with some potential, in $N$ spatial dimensions. It is described by position and momentum operators which satisfy canonical commutation relations of the form
$[Q_{i},Q_{j}]\ =\ 0\,,\qquad[P_{i},P_{j}]\ =\ 0\,,\qquad[P_{i},Q_{j}]\ =\ C\delta_{ij}\,.$ (141)
The elements $\left\{P_{1},\ldots,P_{N},Q_{1},\ldots Q_{N},C\right\}$, where the central element $C$ will be chosen to be i-i, form the Heisenberg Lie algebra $\mathfrak{h}_{2N+1}$. The algebra elements can be identified with $(N+2)\times(N+2)$ nilpotent matrices. For example, for $N=1$, we have
@@
so a general element given by h=aP+bQ+cCh=a\,P+b\,Q+c\,C is identified with
@@
When exponentiated, $Q$ acts on the wavefunctions by multiplying by a coordinate-dependent phase and $P$ generates a translation. This gives rise to the Schrödinger representation of the Heisenberg group. Considering $N$ dimensions, a general element $(\mathbf{a},\mathbf{b},c)\in H_{N+2}$ acts on the wavefunction $\psi(\textbf{q})\in L^{2}(\bf q)$ by
@@
$\bigg[\sum_{i}\left(a_{i}P_{i}+b_{i}Q_{i}\right),\sum_{j}\left(a^{\prime}_{j}P_{j}+b^{\prime}_{j}Q_{j}\right)\bigg]=-iS((\mathbf{a},\mathbf{b}),(\mathbf{a}^{\prime},\mathbf{b}^{\prime})).$ (151)
@@
and satisfy $a_{i},a_{j}^{\dagger}]\ =\ \delta_{ij}$. With these we can define the infinite dimensional oscillator representation $\omega$, which is the minimal faithful representation of the metaplectic group. The oscillator representation is
@@ A.1 Finite fields
The finite-field Heisenberg group also obeys a version of the Stone–von Neumann theorem. For any group $G$ and representation $\rho$, the restriction of the representation to the center $\rho(Z(G))$ must be one-dimensional by Schur’s lemma. Therefore in an irreducible representation, $z\in Z(G)$ acts by multiplication by a (complex) scalar $\chi(z)$. The function $\chi$ is called the central character of $\rho$. In the case of the Heisenberg group, the center is $(\mathbf{0},\mathbf{0},c)\simeq{\mathbb{F}}_{p}$. The statement of uniqueness is that the representations of the Heisenberg group are uniquely (up to unitary equivalences) determined by the value of the central elements in the representation, i.e. by the choice of central character. There are $p$ central characters $\chi(z)=\xi^{z}$, where $\xi$ is a $p^{\text{th}}$ root of unity.
@@
Now $U_{g}$ forms a representation of the symplectic group with dimension $p^{N}$– this is known as the oscillator representation. It is naturally a projective representation, but (unlike in the case for local fields) the phases can be chosen so that it is a genuine (non-projective) representation. We make no distinction between the two (projective / genuine) representations in our notation, though they differ by a phase in the definition of $U_{g}$.
@@
The space of dimension $p^{N}$ (here, the space of functions) which carries a representation of the Heisenberg and symplectic groups is what we call the oscillator representation, denoted $\mathcal{X}$ in our paper. For us, it is spanned by the set $\psi_{c_{1},\ldots,c_{g}}$ for codewords $c_{i}$ of length $2n$,
@@
The oscillator representation $\mathcal{X}$ is an irrep of the Heisenberg group, but actually breaks into two irreps of the symplectic group. On the space of functions, these representations are the even and odd functions, which are preserved by the symplectic group. These subspaces have dimensions $(p^{N}+1)/2$ and $(p^{N}-1)/2$, respectively and correspond to charge-conjugation even (odd). We can see this at work in the case where $p=3$, $n=1$, $g=1$, presented in section 3.3. There we find that the oscillator representation is $3^{2}=9$ dimensional, and it breaks into symplectic irreps of dimension $5$ and $4$. More generally, the group $Sp(2N,{\mathbb{F}}_{q})$ has, apart from the trivial representation, four “minimal” representations with dimensions Gurevich2016
@@ Appendix B Classical groups over finite fields
where $\epsilon=\pm$ is the Witt index of the form $\eta$. We are interested in, c.f. (20),
@@ Appendix C Phase ambiguity for $\boldsymbol{p=3}$ (mod $\boldsymbol{4}$)
where $f(x)\in L^{2}({\mathbb{F}}_{p}^{2ng})$. Now while the first of these expressions has a more natural action, it is the second that is compatible with the oscillator representation being a genuine representation of $Sp(4ng,{\mathbb{F}}_{p})$, see Neuhauser2002 . In other words, if we have $G\times H=O^{+}(2n,{\mathbb{F}}_{p})\times Sp(2g,{\mathbb{F}}_{p})$ as a subgroup of $\mathcal{S}=Sp(4gn,{\mathbb{F}}_{p})$, and $\omega$ is the oscillator representation, we then have the restrictions
@@
This extra phase factor $(\det\mathsf{g})^{\frac{p-1}{2}g}$ (the tensor product with a power of the determinant rep) is sometimes not made explicit in the literature, see however equation (1.4) of Ma2022 . [^37] We thank Shu-Yen Pan for correspondence. It also introduces an ambiguity in the nomenclature of the singlet representation: the case we need for our proof in the main text is where $\pi$ acts trivially, $\pi(\mathsf{g})=\mathbf{1}$; for odd $\frac{p-1}{2}g$ this would correspond to the case where $U_{\mathsf{g}}$ acts as the determinant representation $U_{\mathsf{g}}=\det(\mathsf{g})\mathbf{1}$.